### v1

# Stable Trajectory Clustering: An Efficient Split and Merge Algorithm

Abstract
Clustering algorithms group data points by characteristics to identify patterns. Over the past two decades, researchers have extended these methods to analyze trajectories of humans, animals, and vehicles, studying their behavior and movement across applications. This paper presents whole-trajectory clustering and sub-trajectory clustering algorithms based on DBSCAN line segment clustering, which encompasses two key events: split and merge of line segments. The events are employed by object movement history and the average Euclidean distance between line segments. In this framework, whole-trajectory clustering considers entire entities’ trajectories, whereas sub-trajectory clustering employs a sliding window model to identify similar sub-trajectories. Many existing trajectory clustering algorithms respond to temporary anomalies in data by splitting trajectories, which often obscures otherwise consistent clustering patterns and leads to less reliable insights. We introduce the stable trajectory clustering algorithm, which leverages the mean absolute deviation concept to demonstrate that selective omission of transient deviations not only preserves the integrity of clusters but also improves their stability and interpretability. We run all proposed algorithms on real trajectory datasets to illustrate their effectiveness and sensitivity to parameter variations.
Index Terms:
Trajectory clustering, Whole-trajectory clustering, Sub-trajectory clustering, Stability, DBSCAN
I Introduction
Trajectory clustering is one of the most important data analysis methods developed for analyzing moving objects like humans, animals, and vehicles in various fields and applications. It can be used to identify common behaviors and recognize similar movement patterns, as well as to detect irregular or anomalous movements in different types of trajectory data. Among the use cases of trajectory clustering, we can mention exploring urban travel patterns that reflect the regularity of mobility patterns, which is an important factor in evaluating the optimality of city spatial structure and road network connectivity [1], classifying animal behaviors (foraging, resting, walking) using high-frequency tracking data to study their environmental interactions [2] and clustering animals by species based on their daily movement patterns, which is complicated by the erratic nature of these movements [3]. Trajectory clustering is also used in air traffic categorization [4, 5], traffic network identification [6] and, weather prediction [7].
Anomalies frequently occur in trajectory clustering due to irregularities or deviations in movement patterns. To address this, numerous algorithms have been developed, each designed to effectively detect and analyze these anomalies [8, 9]. Missing, however, is an algorithm that can filter out and ignore insignificant anomalies in trajectories to reveal clustering patterns that are otherwise stable. Researchers have focused in recent years on stable algorithms for clustering data points that select the optimal number of clusters [10]. Said another way, current stable clustering algorithms try to split and merge data points correctly in a way that is robust against random oscillations. Inspired by these developments for point clustering, we propose a stable algorithm for clustering trajectories based on the results of line segment clustering in each time interval. For line segment clustering in a time interval, we use the results from the previous time interval, which reflect the states of moving objects and their changes caused by split and merge events in each time interval. We also present an innovative approach that detects which trajectories can be assigned to their respective cluster despite anomalies, which is a prominent issue in real trajectory data that previous algorithms have not addressed.
Building upon existing trajectory clustering algorithms, we propose algorithms for both whole-trajectory clustering and sub-trajectory clustering. A trajectory can be described as a piecewise linear function, where each segment represents an object’s movement in a specific time interval. Consequently, for trajectory clustering, a set of line segments exists for each time interval, and these sets need to be clustered. For this purpose, Density-Based Spatial Clustering of Applications with Noise (DBSCAN) line segment clustering defines the split and merge procedures of line segments. One key advantage of this approach is that the results from clustering in each time interval can be reused in the subsequent interval, as line segments often retain their previous cluster assignment. In this context, we use the continuous average Euclidean distance to measure the distance between segments. As a result, a data analyst can identify the behaviors of moving objects across all time intervals, determining which objects move together and belong to the same cluster in each interval. The sub-trajectory clustering algorithm focuses on specific time intervals to identify similar sub-trajectories. We employ a sliding window model that utilizes two parameters to recognize temporal patterns and similar sub-sequences. We then present an innovative definition and algorithm for the stable trajectory clustering algorithm, which is a first-of-its-kind advancement. This approach deals with insignificant anomalies and deviations in object movements that obscure broader clustering patterns and can thus be ignored when the goal is quantifying persistent trajectory clusters. It is often important to filter out anomalies in data where imprecision is small relative to the scale of the movement. For instance, in urban traffic analysis, a vehicle making a brief stop at a traffic light or a pedestrian crossing should still be classified within the main traffic cluster. Furthermore, a person briefly deviating from a common trajectory to look at a shop window should not be marked as an outlier in pedestrian movement studies. A related example can be seen in animal movement studies, where a migrating bird or herd of animals temporarily deviates from its migration trajectory to find food or rest, but then returns to their original trajectory afterward. In this algorithm, we employ the mean absolute deviation concept to assess how much each outlier trajectory deviates from its assigned cluster, enabling us to distinguish significant outliers from temporary anomalies. Finally, we apply all algorithms and concepts to two real trajectory datasets and demonstrate the parameter sensitivity and effectiveness of the stable trajectory clustering algorithm based on its impact on data, specifically on how the number of clusters changes after the algorithm is executed. In summary, our contributions are illustrated as a graphical abstract in Fig. 1.
The remainder of this paper is organized as follows: Section 2 covers the concepts and related algorithms for data point clustering, sub-trajectory clustering, whole-trajectory clustering, and stable data point clustering. Section 3 covers the foundational concepts and background information necessary for understanding our work, and then formally defines the problems to be solved. Section 4 addresses how whole-trajectories and sub-trajectories are clustered using the split and merge approaches. Section 5 introduces the principles and definitions of the stable trajectory clustering algorithm. Section 6 evaluates the algorithms and discusses their efficiency. Finally, Section 7 concludes the paper with remarks and directions for future work.
Figure 1: Graphical abstract of the proposed approaches.
II Related Work
In the last two decades, many data point clustering algorithms have been discussed in various fields and applications such as wireless sensor networks [11], flight anomaly detection [12], and image processing [13]. These algorithms try to group similar data points based on their similarities or dissimilarities [14]. There are several types of methods to cluster data points, such as partitioning methods [15, 16, 17], hierarchical methods [18, 19], grid-based methods [20, 21], and density-based methods [22, 23]. Some of these methods are used for clustering data points as well as for trajectory clustering. For example, DBSCAN [22] is a density-based algorithm that is used for clustering both data points and trajectories. It can identify the shape and size of data, can handle noise and outliers efficiently, and does not require the number of clusters as a user-defined input. However, DBSCAN cannot identify clusters with different densities and is sensitive to the choice of user input parameters for clustering.
Trajectory data are recorded in different formats by various devices such as Global Positioning System (GPS) trackers, Radio Frequency Identification (RFID) systems, and smartphones. Numerous algorithms have been designed to cluster these different types of trajectory data. They also contribute to improving research, innovation, and theoretical understanding in various fields and applications. Yu et al. [24] suggested a density-based clustering algorithm for trajectory data streams. They designed an algorithm to handle moving object trajectory data as an incremental line segment stream and introduced a distance measurement approach to cluster those line segments via DBSCAN line segment clustering at the end of the current time interval. Their algorithm also used a structure for updating line segment clusters online. Da et al. [25] dealt with the problem of identifying sub-trajectory clusters at every time interval. To this aim, they defined a structure called a micro-group to take and maintain small groups of moving objects at each time window and presented an algorithm for maintaining each micro-group incrementally to discover patterns. Thus, they suggested a density-based sub-trajectory cluster identification approach by merging micro-groups. Trajectory clustering can be divided into point-based, sub-trajectory-based, and whole-trajectory-based methods. Point-based trajectory clustering is a method of clustering trajectories by comparing specific locations (points) along each trajectory. Instead of looking at the entire trajectory, it focuses on the individual points where the object was at different times, and groups similar ones together. Li et al. [26] offered a group discovery framework to facilitate the efficient online discovery of moving objects traveling collectively. Their framework employed a sampling-independent approach while also naturally supporting the utilization of approximate trajectories. Jensen et al. [27] investigated effectively managing a collection of constantly moving data points in a two-dimensional Euclidean space to maintain a cluster of dynamic data points. Their proposed approach can cluster moving objects incrementally. Tang et al. [28] introduced the smart-and-close discovery algorithm, which efficiently generates companions from trajectory data. In their approach, the objective of the proposed traveling buddy model was to improve the clustering and intersection processes involved in discovering companions. Li et al. [29] recommended utilizing cluster analysis for mobile objects and introduced the notion of moving micro-clusters, which can effectively identify patterns in the mobility of objects even in massive datasets. Gao et al. [30] addressed the challenges of clustering real-time trajectory data generated by GPS-equipped devices. They proposed a novel framework that dynamically models trajectory streams using a self-adaptive k-segment method, considering both spatial and temporal dimensions. This framework supported distributed indexing and processing to handle large-scale data efficiently. It utilized incremental DBSCAN clustering for real-time processing and integrated a two-layered distributed index to enhance scalability. The study’s key contributions included effective trajectory stream modeling, efficient and incremental clustering algorithms, and scalable distributed processing techniques.
Sub-trajectory clustering involves breaking down a trajectory into smaller line segments and clustering them based on similarity in terms of spatiotemporal properties. It can be useful when it is not necessary to analyze the entire trajectory. In contrast, whole-trajectory clustering is more commonly used for identifying patterns over a longer period of time. Lee et al. [31] introduced a novel partition-and-group framework for clustering sub-trajectories, which involves dividing trajectories into a collection of line segments and subsequently grouping similar line segments into a cluster [32]. Buchin et al. [33] presented a linear-time algorithm for calculating the most similar sub-trajectories under equal starting times, while for unequal starting times, they offered $(1+\varepsilon)$-approximation algorithms that were not as efficient. This approach employed a distance measure defined as the average Euclidean distance at corresponding times. Mao et al. [34] used a sliding window model to propose an online algorithm for clustering streaming trajectories. This model contained a micro-clustering that represented clustering trajectory line segments in a current window and a macro-clustering based on micro-clustering over a specified time interval. They defined two data structures for tracking the latest alterations of a cluster trajectory data stream in real-time. Buchin et al. [35] also examined the problem of identifying commuting patterns in trajectories by searching for comparable sub-trajectories. They proposed multiple approximation algorithms and demonstrated that computing and approximating the longest sub-trajectory cluster was as hard as the Max-Clique problem. Liu et al. [36] defined a method for determining the distance between sub-trajectories considering time, space, and direction. They proposed a density trajectory cluster algorithm in which the cluster radius is calculated by the density of the data distribution. Schreck et al. [37] suggested a complete visual-interactive monitoring and control framework for trajectory data analysis based upon an enhanced version of the Self-Organizing Maps algorithm. Their framework allowed users to observe clustering progression visually and manage the algorithm at any desired level. Yang et al. [38] designed a trajectory clustering algorithm based on spatio-temporal density analysis of data and defined two new metrics-NMAST density function and NT factor- in this algorithm. Binh Han et al. [39] proposed a systematic method for whole-trajectory clustering of mobile objects in road networks through their framework TRACEMOB. This includes a unique distance measure that captures the complex characteristics of trajectories, an algorithm to transform these trajectories into multidimensional data points in Euclidean space, and a validation method for assessing clustering quality. The validation approach evaluates the quality of clustering in both the transformed metric space and the original road network space, ensuring comprehensive analysis. Wang et al. [40] designed a grid-based whole-trajectory clustering algorithm that converts the trajectories in the road network space into corresponding grid sequences within grid space. Their approach could determine a series of clusters as well as some abnormal trajectories and GPS points.
Anomaly detection in trajectories involves identifying patterns and behaviors that deviate significantly from the norm [41, 9, 42]. Though the deviations are big enough to qualify as statistical outliers, they are ignorable with respect to the inference on the overall clustering pattern. This is what we address in our work via the stable trajectory clustering algorithm, where the trajectories remain in their respective clusters to the extent possible despite minor anomalies in a few time intervals. However, in previous studies, stable clustering algorithms refer to methods that create and maintain clusters of data points to minimize changes in cluster membership over time. These algorithms ensure efficient communication and resource utilization while considering factors such as distance and connectivity. Shahwani et al. [43] proposed a novel clustering algorithm for Vehicular Ad hoc Networks (VANETs) that enhances cluster stability by utilizing vehicle trajectories, specifically incorporating the traffic regularity of buses, into the clustering process. The algorithm used the Affinity Propagation technique, which traditionally relies on distance metrics, but modified it to also consider the shared trajectories of vehicles. By doing so, the algorithm selects Cluster Heads (CHs), which are central nodes in VANETs responsible for coordinating communication within a cluster of vehicles. These CHs not only have minimal distance from other vehicles but also share longer common trajectories with them, leading to more stable clusters. Simulation results demonstrated that this trajectory-based approach significantly improved cluster lifetime and reduced the frequency of cluster head changes compared to conventional methods, thereby enhancing the overall efficiency and reliability of VANET communications. Tseng et al. [44] suggested a stable clustering algorithm called the Clustering Algorithm using the Traffic Regularity of Buses for VANETs in urban scenarios. The algorithm leveraged the fixed routes and regular schedules of buses to improve cluster stability. It considered vehicle mobility characteristics such as velocity, position, and direction, along with the spatial and temporal dependencies of traffic patterns in urban areas. CATRB used the concept of an equilateral triangle’s centroid, incenter, and circumcenter to select the most appropriate CH. This approach aimed to position the CH at the center of the cluster for efficient control and data transmission. These studies just focused on providing stable algorithms for data point clustering. While stable data point clustering is undoubtedly important, it does not fully address the complexities and dynamic nature of trajectory clustering. Our work aims to fill this gap by introducing and rigorously evaluating a stable algorithm designed for trajectory clustering.
III Preliminaries and Definitions
In this section, we will start by defining key concepts such as trajectory, trajectory line segment, and sub-trajectory, which are essential for formally defining both the whole-trajectory clustering and sub-trajectory clustering problems. Then we will introduce the average Euclidean distance and explain the concepts underpinning the DBSCAN line segment clustering approach.
Definition III.1 (Trajectory).
Geometrically, a trajectory is a piecewise linear function of a moving object that is obtained by connecting geographical points that are recorded at different moments $t_{1},t_{2},...,t_{T}$ by a tracking device. So, a trajectory for the $k$-th object is denoted as:
$\pi_{k}=(\overrightarrow{P_{k}^{1}P_{k}^{2}},\overrightarrow{P_{k}^{2}P_{k}^{3}},...,\overrightarrow{P_{k}^{T-1}P_{k}^{T}}),$
such that each $P_{k}^{i}$ is the object’s position at the $i$-th time-stamp (i.e., $P_{k}^{i}=\left(x_{k}\left(t_{i}\right),y_{k}\left(t_{i}\right)\right)$ where $x_{k}\left(t_{i}\right)$ and $y_{k}\left(t_{i}\right)$ are the longitude and latitude respectively).
Definition III.2 (Trajectory line segment).
$\overrightarrow{P_{k}^{i}P_{k}^{i+1}}$ is a trajectory line segment known as a line segment which is an approximate continuous object movement from position $P_{k}^{i}$ to $P{{}_{k}}^{i+1}$ with constant velocity and zero acceleration in a time interval $\left[t_{i},t_{i+1}\right]$.
Definition III.3 (Sub-trajectory).
A sub-trajectory of trajectory $\pi_{k}$ is characterized by the sub-sequence denoted as $\pi_{k}^{\left[i,j\right]}=(\overrightarrow{P_{k}^{i}P_{k}^{i+1}},\overrightarrow{P_{k}^{i+1}P_{k}^{i+2}},...,\overrightarrow{P_{k}^{j-1}P_{k}^{j}})$ where $1\leq i<j\leq T$. This time interval range is represented by $R_{p}$ with $p=\left[i,j\right]$.
  • Problem 1.
    (Whole-trajectory clustering) Given $n$ trajectories as $\Pi=\left\{\pi_{1},\pi_{2},...,\pi_{n}\right\}$, the aim is to identify patterns and similarities among these trajectories and group them based on their spatiotemporal characteristics into clusters $C=\left\{c_{1},c_{2},...,c_{\#cl}\right\}$. To this end, it is important to know the movement of the objects from time $t_{1}$ to $t_{T}$. However, due to changes in the position of objects over a time interval, it is impossible to identify object movements solely based on their positions at specific moments $t_{1},t_{2},...,t_{T}$. Thus, for $n$ trajectories, there are $n$ line segments at any time interval, which should be clustered by an appropriate similarity criterion. The line segment trajectory clustering results at each time interval contribute to the whole-trajectory clustering of $\Pi$.
  • Problem 2.
    (Sub-trajectory clustering) Let $\Pi=\left\{\pi_{1},\pi_{2},...,\pi_{n}\right\}$ be a set of $n$ trajectories. The purpose is to detect patterns and similarities among these trajectories in $T^{\prime}(<T)$ time intervals such as $R_{1},R_{2},...,R_{T^{\prime}}$. The resemblances can be obtained through $C_{p}^{R_{p}}=\left\{c_{s_{1}},c_{s_{2}},...,c_{s_{num}}\right\}$, in which $c_{s_{k}}$ is the clustering of sub-trajectories $\pi_{1}^{R_{p}},\pi_{2}^{R_{p}},...,\pi_{n}^{R_{p}}$ in time interval $R_{p}$. Thus, the set $C_{s}=\left\{C_{1}^{R_{1}},C_{2}^{R_{2}},...,C_{T^{\prime}}^{R_{T^{\prime}}}\right\}$ is obtained as a series of sets of clustered sub-trajectories.
Definition III.4 (Average Euclidean distance of line segments).
This measure should consider the speed and time of the moving object in a line segment. Hence, the dissimilarity between the two line segments $l_{k}$ and $l_{j}$ is defined as:
[…]
$\begin{cases}x_{k}\left(t_{\sigma}\right)=x_{k}\left(t_{i+1}\right)\times t_{\sigma}+x_{k}\left(t_{i}\right)\times\left(1-t_{\sigma}\right),\\ y_{k}\left(t_{\sigma}\right)=y_{k}\left(t_{i+1}\right)\times t_{\sigma}+y_{k}\left(t_{i}\right)\times\left(1-t_{\sigma}\right).\end{cases}$ (2)
Lemma III.1.
The definition (1) satisfies all metric space axioms, including non-negativity, symmetry, and the triangle inequality.
Proof.
To prove the non-negativity of the metric, we expand the definition as follows:
The Euclidean distance between two moving objects $k$ and $j$ at any moment $t_{\sigma}$, is defined as:
$dist\left(P_{k}^{t_{\sigma}},P_{j}^{t_{\sigma}}\right)$ $=\sqrt{\left(x_{k}\left(t_{\sigma}\right)-x_{j}\left(t_{\sigma}\right)\right)^{2}+\left(y_{k}\left(t_{\sigma}\right)-y_{j}\left(t_{\sigma}\right)\right)^{2}}.$ (3)
Substituting the definitions of $x_{k}\left(t_{\sigma}\right)$ and $y_{k}\left(t_{\sigma}\right)$ from formula (2) into Eq. 3, we have:
$dist\left(P_{k}^{t_{\sigma}},P_{j}^{t_{\sigma}}\right)$ $=\sqrt{\left(At_{\sigma}+B\right)^{2}+\left(Ct_{\sigma}+D\right)^{2}}$
$=\sqrt{A^{2}t_{\sigma}^{2}+2ABt_{\sigma}+B^{2}+C^{2}t_{\sigma}^{2}+2CDt_{\sigma}+D^{2}}$
$=\sqrt{\left(A^{2}+C^{2}\right)t_{\sigma}^{2}+\left(2AB+2CD\right)t_{\sigma}+B^{2}+D^{2}},$
where $A=x_{k}\left(t_{i+1}\right)-x_{j}\left(t_{i+1}\right)-x_{k}\left(t_{i}\right)+x_{j}\left(t_{i+1}\right)$, $B=x_{k}\left(t_{\sigma}\right)-x_{j}\left(t_{\sigma}\right)$, $C=y_{k}\left(t_{i+1}\right)-y_{j}\left(t_{i+1}\right)-y_{k}\left(t_{i}\right)+y_{j}\left(t_{i+1}\right)$ and $D=y_{k}\left(t_{\sigma}\right)-y_{j}\left(t_{\sigma}\right)$. Therefore, the (continuous) average Euclidean distance of two line segments, $l_{k}$ (i.e., $l_{k}=\overrightarrow{P_{k}^{i}P_{k}^{i+1}}$), and $l_{j}$ is defined as follows:
[…]
If we set $a=A^{2}+C^{2},\quad b=2AB+2CD,\quad c=B^{2}+D^{2}$, then
[…]
$=\frac{1}{\tau}\evaluated{\left(\frac{(2at_{\sigma}+b)\sqrt{at_{\sigma}^{2}+bt_{\sigma}+c}}{4a}\right)}_{t_{i}}^{t_{i+1}}$
$-\evaluated{\frac{1}{\tau}\left(\frac{(b^{2}-4ac)\ln(2\sqrt{a}\sqrt{at_{\sigma}^{2}+bt_{\sigma}+c}+2at_{\sigma}+b)}{8a^{\frac{3}{2}}}\right)}_{t_{i}}^{t_{i+1}}.$
Since $a>0$ and $b^{2}-4ac<0$, the quadratic function $at^{2}+bt+c$ is positive definite, ensuring the metric is always non-negative. The identity of indiscernibles and symmetry are evident from the definition of the Euclidean norm in the integral. For the triangle inequality, given three line segments $l_{i}$, $l_{j}$, and $l_{k}$, it can be shown straightforwardly as follows:
$d(l_{i},l_{k})\leq d(l_{i},l_{j})+d(l_{j},l_{k}).$
Line segments can be clustered based on density in any time interval. In this paper, we use DBSCAN line segment clustering, which needs two user-predefined parameters, $\varepsilon$ and $MinLns$. The first one is the maximum distance between two line segments, which is used to locate the line segment in the neighborhood of any other line segment. The second one is the minimum number of line segments that need to be clustered together in a region to be considered dense. Assume that $L=\left\{l_{1},l_{2},...,l_{n_{s}}\right\}$ is the set of unclustered line segments. The principles and concepts of DBSCAN clustering for these line segments are defined as follows.
Definition III.5 ($\varepsilon$-neighborhood).
The $\varepsilon$-neighborhood of line segment $l_{i}\in L$ denoted by $N_{\varepsilon}(l_{i})$ and defined as $N_{\varepsilon}(l_{i})=\left\{\forall l_{j}\>|\>dist(l_{i},l_{j})\leq\varepsilon\right\},l_{j}\in L$.
Definition III.6 (Core line segment).
$l_{i}$ is a core line segment if $|N_{\varepsilon}(l_{i})|\geq MinLns$, where $|N_{\varepsilon}(l_{i})|$ is the size of set $N_{\varepsilon}(l_{i})$.
Definition III.7 (Directly density-reachable).
$l_{i}$ is directly density-reachable from a line segment $l_{j}$ if $l_{i}\>\in N_{\varepsilon}(l_{j})$ and $MinLns\leq|N_{\varepsilon}(l_{j})|$.
Definition III.8 (Density-reachable).
A line segment $l_{i}$ is density-reachable from line segment $l_{j}$ if there is a chain of line segments $l_{i},l_{i+1},...,l_{j}$ such that line segment $l_{k}$ is directly density-reachable from line segment $l_{k+1}$ ($i\leq k\leq j-1$).
Definition III.9 (Density-connected).
A line segment $l_{i}$ is density-connected to line segment $l_{j}$ if there is a line segment $l_{k}$ such that both $l_{i}$ and $l_{j}$ are density-reachable from $l_{k}$.
Definition III.10 (Borderline segment).
A line segment $l_{i}$ is a borderline segment if it is not a core line segment but is directly density-reachable from a core line segment $l_{j}$.
Definition III.11 (Outlier line segment).
A line segment $l_{i}$ is an outlier line segment if it is neither a core line segment nor a borderline segment.
Definition III.12 (Cluster of line segments).
Assume that $\Gamma$ is a non-empty subset of $L$. For $\Gamma$ to be a cluster, two conditions must be satisfied.
  1. 1.
    (Maximality) For any pair of line segments $l_{i}$ and $l_{j}$, if $l_{i}\in\Gamma$ and $l_{j}$ is density-reachable from $l_{i}$, then $l_{j}\in\Gamma$.
  2. 2.
    (Connectivity) For any pair of line segments $l_{i}$ and $l_{j}$ which are in $\Gamma$, $l_{i}$ is density-connected to $l_{j}$.
The definitions above help clarify the key concepts and terminology relevant to our discussion of the stable trajectory clustering algorithm. The next section will explore how line segments can be clustered in each time interval, taking into account their history of clustering in the previous time interval. We then use the results of line segment clustering in each time interval to cluster whole-trajectories and sub-trajectories.
IV Clustering Scheme
As mentioned in previous clustering studies, the split and merge operations are two significant events for moving object clustering  [29, 27]. In this paper, we also support this view of clustering line segments in each time interval. The split event for line segments involves identifying which line segments are separated from each other, while the merge event involves combining line segments that are close to each other and have similar properties.
Algorithm 1 outlines this trajectory clustering scheme. Segmentation is performed to simplify the analysis and processing of trajectories. The algorithm uses DBSCAN line segment clustering concepts to cluster these line segments within each time interval by the split and merge events of line segments. It checks each cluster from the previous time interval for splitting or merging. If a cluster is dense, it may be split; if clusters are mergeable, they may be combined. Finally, it utilizes the line segment clustering results for both whole-trajectory and sub-trajectory clustering.
input : $\Pi$, $\varepsilon,MinLns$, $S$ and $W$.
output : Whole-trajectory cluster set $C$ and sub-trajectory cluster set $C_{s}$.
1 $m\leftarrow T-1;$
2 divide each $\pi_{i}\in\Pi$ into $m$ parts;
3 $C_{curr}\leftarrow DBSCAN(L,\varepsilon,MinLns);$
4 $C_{hist}\left\{1\right\}\leftarrow\left\{C_{curr}\right\};$
5 for $t$ from 2 to $m$ do
6       $C_{new}\leftarrow\emptyset;$
7
8      foreach $c_{i}\in C_{curr}$ do
9             $C_{new}$.union$\big{(}|c_{i}|>1\,?\,\text{Split}(c_{i},\varepsilon,MinLns):c_{i}\big{)};$
10      foreach pairs of $c_{i}$ and $c_{j}$ in $C_{new}$ do
11             if $isMergeable(c_{i},c_{j})$ then
[…]
13                   if $C_{mrg}\sim=\emptyset$ then
14                        remove $c_{i}$ and $c_{j}$;
15                         $C_{new}${end+1}$\leftarrow C_{mrg}$;
16      $C_{curr}\leftarrow C_{new};$
17       $C_{hist}\left\{t\right\}\leftarrow\left\{C_{curr}\right\};$
[…]
19 $C_{s}\leftarrow$ subTraj($C_{hist},m,\Pi,S,W$);
return $C,\>C_{s};$
Algorithm 1 Trajectory clustering
The details of this algorithm are as follows: Firstly, the preprocessing should be run on data trajectories to divide each trajectory into $m$ parts (lines 1 and 2). Then, in the first time interval, for obtaining the first clusters, DBSCAN line segment clustering should be executed (line 3). From the second time interval to the last time interval (lines 5-16), if any cluster of line segments exists, the Split function (Algorithm IV-A) is called (line 8). Two outputs are obtained: either the segments remain in their cluster or they are split from each other. The Merge function (Algorithm 3) should then be executed based on these results (lines 9-15). Eventually, the results of line segment clustering are utilized as inputs for the whole-trajectory clustering (Algorithm 4) and the sub-trajectory clustering (Algorithm 5) algorithms (lines 17 and 18).
IV-A Split of Line Segments
According to definition III.12, let $C_{s}$ be a cluster containing $n_{s}$ line segments in the time interval $\left[t_{i-1},t_{i}\right]$, where $n_{s}\leq n$. In the following, we will provide a definition and explanations regarding the split of the cluster.
Definition IV.1 (Split event).
Cluster $C_{s}$ will be split into at least two non-empty sub-clusters $c_{s_{1}}$ and $c_{s_{2}}$ in the next time interval $\left[t_{i},t_{i+1}\right]$ if no line segment in $c_{s_{1}}$ is density-reachable from any line segment in $c_{s_{2}}$. In other words, if we have two sub-clusters $c_{s_{1}}$ and $c_{s_{2}}$, and for every line segment $l_{i}$ in $c_{s_{1}}$ and every line segment $l_{j}$ in $c_{s_{2}}$, there is no chain of density-reachable line segments connecting $l_{i}$ to $l_{j}$, then $l_{i}$ and $l_{j}$ are not in the same cluster. Therefore, $c_{s_{1}}$ and $c_{s_{2}}$ are considered split sub-clusters.
Suppose that $UC_{s}$ is a set of line segments in the time interval $\left[t_{i},t_{i+1}\right]$ that have previously been in cluster $C_{s}$. To detect whether the cluster $C_{s}$ will be split in the next time interval or not, a random line segment in $UC_{s}$ should be selected initially and then its neighboring line segments are determined based on their average Euclidean distance as sub-set $c_{s_{1}}$. The same process is repeated for the remaining line segments in $UC_{s}$ until all line segments are examined, and no line segments are remaining. So, there is a set of sub-sets $NC_{s}=\left\{c_{s_{1}},c_{s_{2}},...,c_{s_{num}}\right\}$ and we need to determine whether the members of each sub-set can be considered neighbors of each other or not. If there is no chain of line segments that connects two sub-sets, they are considered split from each other; otherwise, they are connected. So, the $NC_{s}$ should be updated such that its members are the new sub-sets where, according to definition III.12, each sub-set satisfies the Maximality criterion. However, for each subset to be a cluster, we also need to check for Connectivity. Consequently, we have sub-sets that are either considered as a sufficiently dense cluster or are recognized only as a low-density cluster of $NC_{s}$ due to violating the Connectivity criterion. It is noteworthy that it is possible to preprocess the line segments before using formula (III.4), meaning we can measure only the Euclidean distance between the start and end points of the line segments without using the average Euclidean distance of line segments. Therefore, the following cases occur between the start and end points of each pair of line segments in $UC_{s}$ at the time interval $\left[t_{i},t_{i+1}\right]$:
  • If the Euclidean distance of two start points and two endpoints is less than or equal to $\varepsilon$, both line segments are in the same sub-set and it is not necessary to calculate the average Euclidean distance of line segments.
  • If the Euclidean distance of start points, endpoints, or both of them is more than $\varepsilon$, calculating the average Euclidean distance of line segments is necessary.
Algorithm IV-A provides the pseudocode for the split event used in trajectory clustering. The algorithm takes a set of densely clustered line segments in a time interval and determines whether these line segments remain connected or are split from each other in the next time interval. If they are split, the resulting sub-clusters are then evaluated to check if they are dense enough. This involves checking the distances between line segments to identify neighbors, ensuring maximal connectivity between clusters, and assessing the density of the clusters based on the presence of core line segments.
input : Line segments set $UC_{s}=\left\{l_{1},l_{2},...,l_{n_{s}}\right\}$ in time interval $\left[t_{i-1},t_{i}\right]$, $\varepsilon$, $MinLns$.
output : New non-empty sub-clusters $NC_{s}=\left\{c_{s_{1}},c_{s_{2}},...,c_{s_{ncl}}\right\}$ in time interval $\left[t_{i},t_{i+1}\right]$.
1 // Neighborhood analysis
2 foreach line segment $l_{i}\in UC_{s}$ do
3      $c_{s_{i}}\leftarrow\left\{l_{i}\right\}$;
4       foreach line segment $l_{j}\in UC_{s}$ do
5             if $dist(l_{i},l_{j})\leq\varepsilon$ then
6                  $c_{s_{i}}$.add$(l_{j})$;
7
[…]
9
10// Maximality
11 foreach pair of $c_{s_{i}}$ and $c_{s_{j}}$ in $NC_{s}$  do
12       separated-clusters $\leftarrow$ true;
13       foreach $l_{s_{i}}\in c_{s_{i}}$  do
14             foreach $l_{s_{j}}\in c_{s_{j}}$  do
15                   if $dist(l_{s_{i}},l_{s_{j}})\leq\varepsilon$ then
16                         separated-clusters $\leftarrow$ false;
17                         $c_{s_{i}}$ and $c_{s_{j}}$ are connected;
18                         goto end-loop;
19
20      end-loop:
21       if separated-clusters is true then
22            $c_{s_{i}}$ and $c_{s_{j}}$ are split;
23// Connectivity
24 foreach $c_{s_{i}}\in NC_{s}$ do
25      if – then
26            absent
27      else
28            c
29            si contains at least one core line segment
30      $c_{s_{i}}$ is a dense enough cluster;
31      $c_{s_{i}}$ is a outlier or low-density cluster;
Algorithm 2 Split
return $NC_{s}$  The detailed steps of this algorithm are as follows: the algorithm finds the neighboring line segment for each line segment in $UC_{s}$ and creates clusters without determining whether they are sufficiently dense or not (lines 2-7). The algorithm then checks the Maximality (lines 9-19) and Connectivity (lines 21-25) criteria, as mentioned in definition III.12.
IV-B Merge of Line Segments
Let $c_{s_{1}}$ and $c_{s_{2}}$ be two independent low-density clusters or sufficiently dense clusters of line segments that comprise at most $n_{s_{1}}$ and $n_{s_{2}}$ line segments, respectively, where $n_{s_{1}}+n_{s_{2}}\leq n$. The following definition explains how dense clusters, low-density clusters, or a combination of both can be merged.
Definition IV.2 (Merge event).
Two clusters $c_{s_{1}}$ and $c_{s_{2}}$ are merged if there are at least two line segments $l_{i}\in c_{s_{1}}$ and $l_{j}\in c_{s_{2}}$ such that the average Euclidean distance of $l_{i}$ and $l_{j}$ is less than or equal to $\varepsilon$.
The following scenarios may occur during the merge event.
  • Merging outlier line segments or low-density clusters.
  • Merging outlier line segments or low-density clusters with dense clusters.
  • Merging dense clusters with each other.
To identify the two line segments that lead to the merge of two clusters, a simple search is conducted between the elements of the two clusters. The search stops as soon as two line segments are found that have an average Euclidean distance of less than or equal to $\varepsilon$. The merge event may decrease the number of clusters, however, and determining whether new clusters have the required density depends on the user input parameters set for the DBSCAN clustering algorithm. Algorithm 3 merges two clusters of line segments based on a proximity threshold and evaluates whether the resulting cluster is dense. The density is determined by the presence of core line segments and the total number of line segments in the merged cluster.
input : Two non-empty low-density or dense enough clusters $c_{m_{1}}$ and $c_{m_{2}}$, $\varepsilon$, $MinLns$.
output : Cluster $c_{m_{1}m_{2}}$.
1 foreach line segment $l_{i}$ in $c_{m_{1}}$ do foreach line segment $l_{j}$ in $c_{m_{2}}$ do if $dist(l_{i},l_{j})\leq\varepsilon$ then $c_{m_{1}m_{2}}\leftarrow c_{m_{1}}\cup c_{m_{2}}$;
2 goto Connectivity;
3 end if
4end foreach
5end foreach
6// Connectivity
7 if $|c_{m_{1}m_{2}}|=|c_{m_{1}}|+|c_{m_{2}}|$ then if $c_{m_{1}m_{2}}$ contains at least one core line segment  then $c_{m_{1}m_{2}}$ is a dense enough cluster; else $c_{m_{1}m_{2}}$ is a low-density cluster; end if
8return $c_{m_{1}m_{2}};$ else return $\emptyset$;
9 end if
Algorithm 3 Merge
In this algorithm, two sufficiently dense or two non-empty low-density clusters are given as inputs. Then it is checked which two line segments can connect two clusters (lines 1-6). If two clusters are connected, then the Connectivity should be checked (lines 7-14). It is noteworthy that in some time intervals, both split and merge events occur. The Split event never executes on the outlier line segments. Instead, only the Merge event is applied to them, which alters the role of the line segments in the clusters they join.
IV-C Whole-trajectory Clustering and Sub-trajectory Clustering
In whole-trajectory clustering, the algorithm considers the entire trajectory of an entity. In other words, features like average Euclidean distance and constant velocity of moving objects in each time interval hold significance for clustering. With this in mind, we can use the clustered line segment information obtained by the split and merge events which we discussed in previous sections. Thus, based on definition III, two whole trajectories $\pi_{i}$ and $\pi_{j}$ can be assigned to the same cluster if their line segments consistently belong to the same cluster; either they are neighbors of each other directly or indirectly across all time intervals (definitions III.8 and III.7). Algorithm 4 provides whole-trajectory clustering by comparing trajectories at multiple time intervals to identify similarities. It counts the number of time intervals during which the line segments of each pair of trajectories are in the same cluster. If this count matches the total number of time intervals, the trajectories in the pair are considered similar because they consistently follow similar paths over time. The algorithm constructs a joint or disjoint graph that represents these similar relationships between trajectories and returns each connected component in this graph as a cluster of similar trajectories.
input : $C_{hist}$, $m$, $\Pi$.
output : $C=\left\{c_{1},c_{2},...,c_{\#cl}\right\}$.
1 $\mathcal{A}_{\mathtt{traj}}\leftarrow\mathbb{O}_{m\times m};$
2 foreach trajectory $\pi_{i}\in\Pi$  do $counter\leftarrow 0;$
3 foreach trajectory $\pi_{j}\in\Pi$ where $j>i$ do for t from 1 to $m$ do for k from 1 to $|C_{hist}\{t\}|$ do if $l_{\pi_{i}}$ and $l_{\pi_{j}}\in C_{hist}\{t\}\{k\}$  then $counter\leftarrow counter+1;$ end if
4end for
5end for
[…]
7end foreach
[…]
$\textnormal{{return}}\;C;$
Algorithm 4 Whole-trajectory clusteirng
In this algorithm, a matrix is used to determine which whole trajectories are in the same cluster in all time intervals (lines 4-9). Subsequently, the graph structure is employed to determine the whole-trajectory clusters. Each graph represents a cluster of whole trajectories. Besides that, a node corresponds to a trajectory, and the edge between two nodes signifies their neighborhood (line 10). For sub-trajectory clustering, we use the sliding window model. This model is appropriate when we aim to detect specific ranges within a given dataset and assess the patterns, trends, or deviations within those ranges  [45, 46, 47]. As mentioned in the definition III, the goal of sub-trajectory clustering is to recognize the time interval ranges like $R_{p}$ and cluster the sub-trajectories in this range as $C_{p}^{R_{p}}$. Utilizing the sliding window model as a tool can facilitate the process of finding these time interval ranges more easily. Therefore, we should first set the size of the sliding window and identify how many steps we want to move it. We consider two parameters $W$ and $S$ as the sliding window size and the movement step forward, respectively, which depend on the time interval range. To give an illustration, if each time interval range is considered 5 seconds, for $W=3$ and $S=1$, the sliding window encompasses the data of 3 consecutive time intervals, which are 15 seconds, and moves forward 5 seconds for each step (Fig. 2(e)). The window moves across all time intervals and according to the window size, contains the line segment clusters obtained in the previous section by the split and the merge events. In the current window, the consecutive time intervals that incorporate similar line segment clusters are determined as sub-trajectory clusters. Based on the parameter $S$, moving the sliding window either combines similar sub-trajectory clusters with the previous ones or creates new sub-trajectory clusters. If $S<W$, then the consecutive windows contain common time intervals and line segment clusters. Algorithm 5 describes sub-trajectory clustering. It processes pre-clustered line segments to identify and cluster similar sub-trajectories. It employs a sliding window approach, iterating over the line segments according to parameters $S$ and $W$. For each window, sub-trajectories are extracted and clustered, with the results saved incrementally. After processing all windows, consecutive clustering results are compared to identify similar and different ranges. Continuous ranges of similar results are merged, while different ranges are noted separately. The algorithm then compiles these identified ranges into a final list of clustered sub-trajectories, resulting in a set of clusters that represent similar segments within the overall trajectory data.
input : $C_{hist}$, $m$, $S$ and $W$.
output : $C_{s}=\left\{C_{1}^{R_{1}},C_{2}^{R_{2}},...,C_{m}^{R_{m}}\right\}$.
1 for i from 1 to $|C_{hist}|$+W-1 by step S do $W_{s}\leftarrow i;$ $W_{e}\leftarrow i+W-1;$
[…]
4 $p\leftarrow[W_{s},W_{e}];$
5 $C_{i}^{R_{p}}\leftarrow$ wholeTraj$(C_{new},R_{p},\Pi_{new})$;
6 $W_{temp}\{i\}\leftarrow C_{i}^{R_{i}};$
7 end for
[…]
9 for j from 1 to $W_{\text{move}}-1$ do $\mathcal{B}_{\text{sim}}\leftarrow(W_{\text{temp}}\{j\}=W_{\text{temp}}\{j+1\})\;?\;1:0$;
10 end for
[…]
12 foreach $i\in\{i^{\prime}\mid\mathcal{B}_{\text{sim}}[i^{\prime}]=0\}$ do $W_{s}\leftarrow(i-1)\times S+1;$ $W_{e}\leftarrow W_{s}+W;$
13 $p\leftarrow[W_{s},W_{e}];$
14 $C_{s}.\text{add}(C_{p}^{R_{p}});$ end foreach
[…]
16 for i from 1 to $|\mathcal{I}_{\text{sim}}|$ do $e\leftarrow$lastElement$(\mathcal{I}_{\text{sim}}\{i\})$;
[…]
18 $W_{e}\leftarrow((e-1)\times S+1)+W;$
19 $p\leftarrow[W_{s},W_{e}];$
20 $C_{s}$.add$(C_{p}^{R_{p}});$ end for
21
return $C_{s};$
Algorithm 5 Sub-trajectory clusteirng
The detailed steps of this algorithm are as follows: Firstly, based on the parameters $S$ and $W$, the trajectories within each window are clustered (lines 1-8). Then, each pair of continuous windows should be compared (lines 10 and 11). Continuous ranges with identical results are therefore merged and identified as a single range in which the sub-trajectories all have the same patterns (lines 13-23). Both whole-trajectory and sub-trajectory clustering follow the same procedures, utilizing the results of the split and merge events of line segment clustering in each time interval. Nevertheless, while the whole-trajectory clustering algorithm focuses on patterns and similarities across entire trajectories, sub-trajectory clustering examines specific parts of trajectories via a sliding window.
V Stable Trajectory Clustering Algorithm
In this section, we address the concept of stability and how it can be applied in the proposed trajectory clustering algorithms.
Definition V.1 (Stable trajectory clustering).
To further enhance the results of trajectory clustering, we can evaluate the outlier trajectories, which are the trajectories of moving objects that deviate from their respective clusters for a few intervals and create anomalies. Such anomalies may be caused by noise in GPS trackers [8] or conditions that are dependent on the nature of the object’s movement. Sometimes minor deviations are often inconsequential and can be disregarded. Whether or not deviations should be ignored depends on the number of time intervals over which the objects’ trajectories remain split from their clusters. Additionally, the magnitude of their deviation from the clusters within these time intervals is important. Indeed, we aim to determine whether, despite the anomaly causing a trajectory to split from its cluster temporarily, it still belongs to the same cluster, or if this anomaly represents a genuine split. As an example, the trajectories of six moving objects from top to bottom in different colors are shown in Fig. 2(a). There are two vertical and horizontal crossed corridors in this example. In the middle, due to the anomalies, the sub-trajectory clustering and whole-trajectory clustering have detected five clusters and three clusters, respectively (Fig. 2(b) and Fig. 2(c)). Now, if the objects spend significantly more time traversing the vertical corridor compared to the horizontal one, the clustering results may be different. In fact, the stability of the trajectory can smooth the impact of movements in the horizontal corridor, and different results may be achieved as shown in Fig. 2(d) and Fig. 2(e), respectively.
Figure 2: Trajectory clustering based on the concept of stability. (a) The trajectories of six moving objects from top to bottom. (b) Detection of five clusters by sub-trajectory clustering. (c) Identification of three clusters by whole-trajectory clustering. (d) Disregarding the mid-trajectory anomalies and considering three sub-trajectories within a single cluster. (e) Ignoring the mid-trajectory anomalies completely and considering all trajectories as belonging to one cluster.
Building upon the discussions of whole-trajectory clustering and sub-trajectory clustering presented earlier, we can now obtain trajectory clusters and outliers. Then, the outlier or split trajectories can be assessed to determine whether their split should be considered a real split or not. Suppose $C_{num}$ represents a cluster of trajectories and $\pi_{a}$, $\pi_{b}$ are two outlier or split trajectories which, due to the anomalies in some time intervals, do not belong to the cluster $C_{num}$. Thus, they should be evaluated by the stable trajectory clustering algorithm. In the following, we will begin this evaluation with $\pi_{a}$. We intend to determine the distance between the modified outlier trajectory of $\pi_{a}$ that is denoted as $\pi^{\prime}_{a}$ and the members of cluster $C_{num}$ such as $\pi_{c}$ using definition III.4. To this aim, suppose $l_{\pi_{a}}$ represents a line segment of $\pi_{a}$ and $l_{\pi_{c}}$ is a line segment of $\pi_{c}$ in $lc^{R_{i}}$ which is the current line segment cluster in $i$-th time interval. For these two outlier trajectories, two states may occur for their line segments in this time interval:
  • First indicates belonging to the current line segment cluster (e.g. $l_{\pi_{a}}\in lc^{R_{i}}$).
  • Second refers to deviating from the current line segment cluster (e.g. $l_{\pi_{a}}\notin lc^{R_{i}}$).
This process is illustrated in Fig. 3.
Figure 3: Two states for outlier trajectories $\pi_{a}$ and $\pi_{b}$. The trajectories in black are members of cluster $C_{num}$. The dashed line segments represent the approximate movement of objects that have deviated from their respective line segment clusters.
Suppose the first state mentioned occurs within this time interval; in that case, it is sufficient to calculate the average Euclidean distance between $l_{\pi_{a}}$ and $l_{\pi_{c}}$ directly. This is because both line segments are in the same line segment cluster and are neighbors (definitions III.7 to III.9); otherwise, the second state occurs and the amount by which the distance between $l_{\pi_{a}}$ and $lc^{R_{i}}$ needs to be reduced for $l_{\pi_{a}}$ to become a member of $lc^{R_{i}}$ should be determined. Consider $l_{N}$ as a borderline segment (definition III.10) in $lc^{R_{i}}$. It is evident that the average Euclidean distance between $l_{\pi_{a}}$ and its nearest line segment $l_{N}$ exceeds $\varepsilon$ by $\delta_{i}$. This is because the distance being less than or equal to $\varepsilon$ implies that $l_{\pi_{a}}$ and $l_{N}$ are in the same cluster and are neighbors. Here, $\delta_{i}$ is the deviation in the $i$-th time interval. Accordingly, the average Euclidean distance between $l_{\pi_{a}}$ and $l_{\pi_{c}}$ in the current line segment cluster of the $i$-th time interval should be reduced by $\delta_{i}$. Thus, line segment $l_{\pi_{a}}$ is modified as $l^{\prime}_{\pi_{a}}$. This modified line segment shows the most likely approximate movement of object $a$ in this time interval. Therefore, the piecewise function for obtaining the set of average Euclidean distances in $T-1$ time intervals is defined as follows:
$D(\pi^{\prime}_{a},\pi_{c})_{i}^{R_{i}}=\begin{cases}dist(l_{\pi_{a}},l_{\pi_{c}}),&lc^{R_{i}}(l_{\pi_{a}})=lc^{R_{i}}(l_{\pi_{c}}),\\ dist(l_{\pi_{a}},l_{\pi_{c}})-\delta_{i},&o.w,\\ \end{cases}$ (5)
where $i$ takes values in the range from $1$ to $T-1$ and $\delta_{i}=l_{N}-\varepsilon$. This process is shown in Fig. 4. The maximum value of these distances represents their overall average Euclidean distance and is denoted as $M_{(\pi^{\prime}_{a},\pi_{c})}$. It indicates the maximum possible distance at which the outlier trajectory $\pi_{a}$ would be considered a member of cluster $C_{num}$, disregarding its anomalies.
$M_{(\pi^{\prime}_{a},\pi_{c})}=\max\{D(\pi^{\prime}_{a},\pi_{c})_{i}^{R_{i}}\quad|\quad i=1,\ldots,T-1\}.$ (6)
The same processes occur not only for $\pi_{a}$ and the other members of cluster $C_{num}$ but also for $\pi_{b}$ and each member of cluster $C_{num}$. Eventually, the minimum of the maximum values is determined as $\mu_{min}$. It indicates which outlier trajectories $\pi_{a}$ and $\pi_{b}$, disregarding their anomalies, are more similar to and closer to the members of cluster $C_{num}$. Given that $N_{C}$ and $N_{O}$ represent the number of cluster members and outlier trajectories, respectively, $\mu_{min}$ is defined as follows:
$\mu_{min}=\min\left\{\displaystyle M_{(\pi^{\prime}_{i},\pi_{j})}\;\middle|\;i=1,\ldots,N_{O},\;j=1,\ldots,N_{c}\right\}.$ (7)
Figure 4: The average Euclidean distance between $l_{\pi_{a}}$ and $l_{\pi_{c}}$ with the assistance of the nearest line segment $l_{N}$ in current line segment cluster. $l_{\pi_{a}^{\prime}}$ is denoted as the approximate movement of the object in the $i$-th time interval.
As mentioned in definition V.1, the stable trajectory clustering algorithm aims to assess whether it is possible to ignore the anomalies that caused an outlier trajectory split from its respective cluster or not. The number of time intervals and the deviation amount in these intervals are important factors for deciding whether or not to ignore anomalies. Specifically, one of the following four cases may occur:
Case 1.
The outlier trajectory deviates from its cluster in a few time intervals, but the deviation in these time intervals is large.
Case 2.
The outlier trajectory deviates from its cluster in many time intervals, and the deviation in these time intervals is also large.
Case 3.
The outlier trajectory deviates from its cluster in many time intervals, but the deviation in these time intervals is small.
Case 4.
The outlier trajectory deviates from its cluster in a few time intervals, and the deviation in these time intervals is also small.
Given these cases, we can use the mean deviation concept, which is defined as follows:
Definition V.2 (Mean absolute deviation).
The mean absolute deviation is a statistical measure to determine how far the data are situated from the mean. It measures the average absolute difference between each data point and the mean. The formula (8) depicts the mean absolute deviation (MAD) for a set of $n$ data points such as $X=\left\{x_{1},x_{2},...,x_{n}\right\}$, where $x_{i}\in{\rm I\!R}$ and $\mu$ represents their mean:
$MAD=\frac{1}{n}\sum_{i=1}^{n}\absolutevalue{\mu-x_{i}}.$ | . (8)
In this formula, the deviations to the left and right of the mean are important. Thus, the sum of absolute differences between each right deviation (RD), which refers to the deviation of values that are greater than the mean, is denoted as the right mean deviation (RMD), while the sum of absolute differences between each left deviation (LD), which refers to the deviation of values that are less than the mean, is denoted as the left mean deviation (LMD). If LMD and RMD are not equal, then the data distribution is asymmetric or skewed.
Based on the aforementioned concepts and $\mu_{min}$ (defined in Eq. 7), the goal is to determine which two outlier trajectories can belong to the cluster $C_{num}$, despite anomalies in a few time intervals. To put it differently, it is sufficient only to determine which members of cluster $C_{num}$ can be the neighbors of $\pi_{a}$ and $\pi_{b}$. To this aim, for each outlier trajectory like $\pi_{a}$ and each member of cluster $C_{num}$, the real average Euclidean distances of their line segments that are greater than $\mu_{min}$ are considered as RD and the real average Euclidean distances that are less than $\mu_{min}$ are regarded as LD. Thus, the sum of absolute differences between each right deviation and $\mu_{min}$ is denoted as RMD, while the sum of absolute differences between each left deviation and $\mu_{min}$ is denoted as LMD. These definitions are as follows:
$RMD=\sum_{RD>\mu_{min}}\absolutevalue{RD-\mu_{min}},$ (9)
$LMD=\sum_{LD<\mu_{min}}\absolutevalue{LD-\mu_{min}}.$ | .
Taking into account the aforementioned formula, and based on the fact that the average Euclidean distance of an outlier trajectory to cluster $C_{num}$ must be $\mu_{min}$ for it to be considered a member, if $RMD$ is greater than $LMD$, then a real split has occurred, and the deviations cannot be ignored. Specifically, this occurs when the deviation in the time intervals has been very significant (Case 1. or Case 2.) or the number of time intervals in which the trajectory deviates from its cluster is more than the number of time intervals in which the trajectory is in its cluster (Case 3.); in contrast, if RMD is less than LMD, then, despite the anomalies, they belong to the same cluster and are stable in cluster $C_{num}$ (Case 4.). The stable trajectory clustering is presented in Algorithm 6. This method improves trajectory clustering results by evaluating outlier trajectories and their association with existing clusters. It begins by checking if there are any outlier trajectories. If outliers exist, the algorithm iterates through each cluster, comparing each trajectory’s line segments with those of other trajectories, and calculating distances and deviations. It identifies the maximum distance among all these deviations and sets a minimum threshold, $\mu_{min}$. The algorithm further refines the cluster assignments by computing real distances and evaluating right and left deviations, deciding which outlier trajectories can remain in the respective cluster despite their deviations. If any cluster assignments are altered, it updates the clusters and returns the new set; otherwise, it returns the original set.
Input : Cluster set $C$, Outlier trajectory set $\mathcal{O}$, $\varepsilon$.
Output : New cluster set $C$.
1 if $\mathcal{O}=\emptyset$ then return $C$ end if
2foreach cluster $c_{k}\in C$ do $\mathcal{M}_{k}\leftarrow c_{k}.$allMembers;
3 $\mathcal{O}_{k}\leftarrow$ assignOutliers$(c_{k},\mathcal{O})$;
4 $\mu_{\min}\leftarrow\infty$;
5 foreach $\pi_{m}\in\mathcal{M}_{k},\pi_{o}\in\mathcal{O}_{k}$ do $M_{(\pi_{m},\pi_{o})}\leftarrow-\infty$;
6 for $t$ from 1 to $T-1$ do
7if $lc^{R_{t}}(l_{\pi_{m}})=lc^{{R_{t}}}(l_{\pi_{o}})$ then $D(\pi_{m},\pi_{o})_{t}^{R_{t}}\leftarrow\text{dist}(l_{\pi_{m}},l_{\pi_{o}})$; else $\delta\leftarrow\text{nearestDistance}(\pi_{o}^{t},\mathcal{M}_{k}^{t})-\varepsilon$;
8 $D(\pi_{m},\pi_{o})_{t}^{R_{t}}\leftarrow\text{dist}(l_{\pi_{m}},l_{\pi_{o}})-\delta$; end if
[…]
11end foreach
12foreach $\pi_{o}\in\mathcal{O}_{k}$ do foreach $\pi_{m}\in\mathcal{M}_{k}$ do $\text{LMD}\leftarrow 0$, $\text{RMD}\leftarrow 0$;
13 for $t$ from 1 to $T-1$ do $D(\pi_{m},\pi_{o})_{t}^{R_{t}}\leftarrow\text{dist}(l_{\pi_{m}},l_{\pi_{o}})$;
14 if $D(\pi_{m},\pi_{o})_{t}^{R_{t}}<\mu_{\min}$ then $\text{LMD}\leftarrow\text{LMD}+(\mu_{\min}-D(\pi_{m},\pi_{o})_{t}^{R_{t}})$; else $\text{RMD}\leftarrow\text{RMD}+(D(\pi_{m},\pi_{o})_{t}^{R_{t}}-\mu_{\min})$; end if
15end for
[…]
17 break; end if
18end foreach
19
20 end foreach
return $\text{updated}\;C$;
Algorithm 6 Stable trajectory clustering
The details of this algorithm are as follows: Initially, a function is defined to identify clusters and respective outlier trajectories (lines 3 and 4). Afterwards, an iteration is performed for each cluster and its outlier trajectories. This iteration tries to use the mean absolute deviation concept to find the outlier trajectories that, despite their anomalies, still belong to their respective clusters (lines 7- 28). Finally, if the clusters change, they will be updated and returned as a new set (line 29). The stable trajectory clustering algorithm can be utilized to analyze outlier trajectories, which due to their anomalies, are split from their respective cluster in some time intervals. In this analysis, it is important to consider both the number of time intervals in which deviations occur and the magnitudes of those deviations. We therefore used the concept of mean absolute deviation to decide if, despite the anomalies, outlier trajectories could be assigned to their nearest cluster. The stable trajectory clustering algorithm can be applied to both whole-trajectory clustering and sub-trajectory clustering problems.
VI Empirical Results
In this section, we empirically validate our proposed algorithms, which are the whole-trajectory clustering, the sub-trajectory clustering, and the stable trajectory clustering algorithms, using real-world trajectory data and evaluating their runtime performance.
VI-A DataSets and Environment Setup
We use two kinds of real trajectory datasets. Dataset 1 is related to a four-way traffic intersection, including maneuvers such as turns and U-turns, and contains 1900 trajectories [48]. Dataset 2, which is known as the Vehicle Energy Dataset (VED), includes information from 383 privately-owned cars in Ann Arbor, Michigan, USA  [49]. This Dataset contains the trajectories of cars driving on both highways and traffic-dense areas from November 2017 to November 2018 and was collected by onboard OBD-II loggers. All algorithms were implemented in Matlab 2020 and executed on a laptop with an Intel Core(TM) i7-4510U processor and 12GB of memory. Before running the algorithms, the data were preprocessed by removing duplicated spatial coordinates and standardizing all time intervals.
VI-B Effectiveness and Parameter Sensitivity
First, we will consider ranges of values for the two user-specified DBSCAN parameters, $\varepsilon$, and $MinLns$. For each set of DBSCAN parameters, we then apply the whole-trajectory clustering algorithms on Dataset 1. In sub-trajectory clustering, we further explore the effects of the sliding window parameters $S$ and $W$ on Dataset 1. Following that, we execute the stable trajectory clustering algorithm and analyze its effectiveness in Dataset 2. Sensitivity analysis for DBSCAN line segment clustering involves evaluating the number of clusters and cluster density across the ranges of user-specified input parameters. These two parameters can be tuned depending on the characteristics of the data, and as mentioned before, their values impact the clustering result. Specifically, larger values of $\varepsilon$ result in larger neighborhoods, which tend to produce larger and less compact clusters. In contrast, the value of $MinLns$ plays a significant role in assigning core or border status to line segments. Higher values require more neighboring line segments to classify a segment as core, potentially reducing the number of clusters or affecting cluster boundaries. Fig. 5 aims to provide a general overview of the impact of these two parameters on whole-trajectory clustering for Dataset 1. As expected, the number of clusters changes as a function of the values of $\varepsilon$ and $MinLns$. A similar analysis can be conducted to examine the influence of other values for these parameters, as illustrated in Fig. 6.
Figure 5: Effect of $\varepsilon$ and $MinLns$ on the number of clusters.
Figure 6: Whole-trajectory clustering on Dataset 1. Increasing $\varepsilon$ (from 8 to 10) and $MinLns$ (from 5 to 80) leads to denser clusters, alters their count, and impacts the number of outlier trajectories. These changes stem from line segment clustering in each time interval, impacting the overall whole-trajectory clustering.
In sub-trajectory clustering, both the DBSCAN parameters and the sliding window parameters affect the results. The effects of the values $\varepsilon$ and $MinLns$ are the same as those mentioned for whole-trajectory clustering. Thus, we only present the results of varying the sliding window parameters $S$ and $W$.
Figure 7: Sub-trajectory clustering on Dataset 1 during the second iteration of the algorithm. (a) Comparison of sub-trajectory clustering results using different step sizes ($S=3$ and $S=11$) with a fixed window size ($W=15$). (b) Comparison of sub-trajectory clustering results using different window sizes ($W=15$ and $W=21$) with a fixed step size ($S=5$).
Fig. 7 demonstrates the effects of varying $S$ and $W$ on sub-trajectory clustering for Dataset 1, which contains trajectories divided into 50-time intervals. The results correspond to the second iteration of the algorithm. For a fixed window size $W$, increasing the parameter $S$ causes the sliding window to move forward more significantly during the same iterations of the algorithm (7(a)). On the other hand, the larger the sliding window size, the more extended time intervals it encompasses, leading to a wider range of sub-trajectories (7(b)). The execution of the whole-trajectory clustering algorithm on Dataset 2 is illustrated in Fig. 8. For further clarity, magnified views within the figures are provided, highlighting parts of cluster $C_{6}$ and $C_{9}$, along with their outlier trajectories in detail. After applying the stable trajectory clustering algorithm, the minor deviations of the outlier trajectory are disregarded, and the trajectories are reassigned to their cluster (8(b)).
Figure 8: The stable trajectory clustering algorithm results on Dataset 2 over 101-time intervals. (a) The whole-trajectory clustering result before running the stable trajectory clustering algorithm ($\varepsilon=0.006$ and $MinLns=2$). The magnified view of $C_{6}$ and $C_{9}$ includes outlier trajectories that temporarily split from their cluster for a few seconds. (b) The whole trajectory clustering result after running the stable trajectory clustering algorithm. The magnified view shows that former outlier trajectories are added to $C_{6}$ and $C_{9}$.
As discussed in section V, in the stable trajectory clustering algorithm, the number of time intervals and the deviation amount in these intervals are important factors in deciding whether or not to ignore anomalies. Based on these two factors, $\mu_{min}$ is calculated for each of 12 clusters and their corresponding outlier trajectories. For example, this value for cluster $C_{6}$ and its 9 outlier trajectories is $0.0074$. Out of these, 4 are correctly assigned to their respective clusters after applying the stable clustering algorithm. Fig. 9 illustrates the temporal distance of outlier trajectories to their nearest cluster members, recognized as the distance from cluster $C_{6}$. Each box plot represents the distribution of the distances between the segments of an outlier trajectory and the segments of its closest trajectory within the cluster across 101-time intervals. As is clear, outlier trajectories with a mean less than $\mu$ are identified as non-outlier trajectories. This means that the stable trajectory clustering algorithm can ignore insignificant anomalies in a few time intervals. These time intervals are shown in Fig. 10. This figure depicts each outlier trajectory as a series of 101 small rectangles, each corresponding to a time interval. It illustrates not only the time intervals during which each outlier trajectory remained within cluster $C_{6}$ but also those during which it was temporarily split from the cluster.
Figure 9: Temporal distance of outlier trajectories from cluster $C_{6}$. In these box plots, each line represents the average distances over 101-time intervals. The dashed horizontal line indicates $\mu_{min}$, a threshold value calculated to indicate the maximum acceptable average distance for identifying non-outlier trajectories.
Figure 10: Temporal membership of outlier trajectories in cluster $C_{6}$ over 101-time intervals. Blue rectangles represent the time intervals during which the outlier trajectory remains within the cluster, while pink rectangles signify the time intervals when it is outside the cluster. Non-outlier trajectories are represented by the darker blue and darker pink areas, based on indicated non-outlier trajectories in Fig. 9. The darker pink regions are also hatched, indicating time intervals where outlier trajectory deviations were disregarded.
As mentioned before, VED (Dataset 2) provides a rich and diverse collection of real-world driving data, including GPS trajectories and energy consumption records, offering an invaluable resource for analyzing driving behavior and energy use. Although the stable clustering algorithm on the VED data trajectories can detect minor deviations in outlier trajectories, it considers those with sharp changes in direction, which are highly noticeable, as real outlier trajectories. This process improves the quality and reliability of the clustering results, as it considers driving patterns that reflect typical behavior. For example, a trajectory that was mistakenly labeled as an outlier due to a brief stop or a slight turn can now be correctly included in its cluster, improving the representation of real driving behavior. The improved clusters lead to more accurate energy consumption analyses and make the analysis more reliable for real-world applications like eco-driving strategies, trajectory optimization, and traffic flow studies. This filtering slightly reduces the average Silhouette coefficient score. This metric quantifies how well data trajectories fit within their clusters (cohesion) and how distinctly they are separated from others (separation). This reduction is due to assigning some outlier trajectories to their respective clusters after applying the stable trajectory clustering algorithm. The Silhouette score is calculated for Dataset 1 and Dataset 2 using different DBSCAN parameter settings before and after executing the stable trajectory clustering algorithm, as shown in Table I. In this table, in addition to this metric, the standard deviation is also computed to indicate the dispersion of the Silhouette scores of trajectories and the range within which these values vary. The key aspect for comparison is the difference in the Silhouette score before and after executing the algorithm. The smaller this difference, and the more trajectories successfully assigned to their respective clusters, the more effective the algorithm is. A lower difference indicates that the algorithm has successfully identified outlier trajectories with minor deviations, thus preserving the overall clustering structure and maintaining the average Silhouette score with minimal change. It should be noted that an outlier trajectory may either connect at least two clusters or be assigned to a single cluster by the stable trajectory clustering algorithm. This occurs because different segments of the outlier trajectory may belong to different clusters. After running the algorithm, the Silhouette score is evaluated for both cases, and the outlier trajectory is ultimately placed where it results in the smallest decrease in the Silhouette score.
Dataset DBSCAN Before STC After STC
$\varepsilon$ MinLns #cl Out. Silhouette #cl Out. Silhouette
#1 $2$ $0.0075$ $13$ $13$ $0.640\pm 0.19$ $13$ $6$ $0.607\pm 0.20$
$2$ $0.0060$ $12$ $18$ $0.653\pm 0.21$ $12$ $7$ $0.617\pm 0.22$
$3$ $0.0065$ $11$ $17$ $0.661\pm 0.19$ $11$ $9$ $0.623\pm 0.20$
#2 $8$ $5$ $18$ $121$ $0.709\pm 0.19$ $18$ $6$ $0.702\pm 0.19$
$10$ $20$ $15$ $42$ $0.750\pm 0.12$ $15$ 00 $0.747\pm 0.12$
$60$ $80$ $11$ $40$ $0.773\pm 0.08$ $11$ 00 $0.722\pm 0.17$
TABLE I: Comparison of clustering results before and after applying the stable trajectory clustering (STC) algorithm for two datasets. The results show how STC improves the cluster structure as it assigns outlier trajectories (Out.) to their respective clusters, which causes a slight reduction in the Silhouette score. A smaller difference in the Silhouette score before and after STC indicates higher effectiveness.
Table II demonstrates the runtime performance of trajectory clustering approaches, where simpler datasets and lower parameter settings yield faster results, while more complex datasets and higher parameter settings require significantly more computation time.
Dataset Split-Merge Whole-Traj. Sub-Traj.
Cl. Algo. STC Cl. Algo. STC
$\varepsilon$ MinLns Time Time Time Time S W Time
#1 $10$ $20$ $353$ $32$ $737$ $172$ $3$ $15$ $1894$
$20$ $60$ $389$ $35$ $957$ $126$ $5$ $20$ $1423$
$40$ $80$ $321$ $31$ $845$ $158$ $4$ $10$ $1787$
#2 $0.0060$ $2$ $15$ $1$ $15$ $15$ $1$ $10$ $69$
$0.0028$ $3$ $13$ $0.9$ $21$ $19$ $1$ $20$ $211$
$0.0040$ $4$ $14$ $1$ $22$ $25$ $2$ $20$ $130$
TABLE II: Runtime performance comparison of trajectory clustering algorithms: Split-Merge (with $\varepsilon$, MinLns), Whole-Trajectory, and Sub-Trajectory methods with parameters $S$ and $W$. Times reported in seconds for Datasets 1 and 2.
VII Conclusion
Trajectory clustering algorithms can analyze either whole trajectories or subsets of them. In this paper, we proposed both a whole-trajectory clustering algorithm and a sub-trajectory clustering algorithm. To discover the continuous object movements in each time interval, the split and merge procedures of line segments were implemented via DBSCAN line segment clustering. We employed the average Euclidean distance to compute the distance between the line segments. In sub-trajectory clustering, we employed the sliding window model to recognize similar ranges of sub-trajectories. We introduced a novel algorithm, called the stable trajectory clustering algorithm, that represents an advanced leap forward in the literature on trajectory clustering. This algorithm addresses the situation when a moving object splits from its cluster for some time intervals and again joins. Ignoring these minor anomalies on large scales assists analysts in performing trajectory clustering without being distracted by insignificant deviations, maintains the integrity of the data, and improves the precision of clustering results. For instance, vehicles stopping at traffic lights, pedestrians pausing at shop windows, or animals detouring briefly during migration should still be considered part of the main movement cluster. To determine whether the anomalies of outlier trajectories are real or temporary, the number of time intervals in which deviations occur and the magnitudes of those deviations are important factors. Thus, we used the mean absolute deviation concept to analyze these factors. The empirical results on real data illustrated the effectiveness of the algorithm in detecting anomalies. Furthermore, the Silhouette score was used to evaluate the quality of the clustering results before and after applying the algorithm. The comparison showed that the effectiveness and robustness of the algorithm increase when two conditions are met: (1) the difference in Silhouette scores before and after applying the algorithm is minimized, and (2) a greater number of outlier trajectories with minor anomalies are correctly reassigned to their respective clusters. The proposed algorithms exhibit sensitivity to the input parameters $\varepsilon$ and $MinLns$, with optimal values being highly dependent on the characteristics of the moving data. This represents a limitation of our approach and warrants further investigation in future research. Additionally, exploring alternative concepts or methods to assess the stable trajectory clustering, beyond the mean absolute deviation approach used in this study, could provide valuable insights.
Acknowledgment
This work was partially funded by the Center of Advanced Systems Understanding (CASUS), which is financed by Germany’s Federal Ministry of Education and Research (BMBF) and by SMWK with tax funds on the basis of the budget approved by the Saxon State Parliament.